6A. Construct the titration curve for a 20.00 m, mixture solution of 0.0800 M in I and 0.0800M in Cl titrated with 0.1000 M AgNO3. Calculate the pAg values of the titration solution atter addition of 6.00mL, 16.00 mL, and 26.00 mL of 0.1000 M AgNO solution. 18) Kp AgCl = 1.77 × 10-10ー[Ag+][Cl] Ksp Agl 8.3 x 101-AgI
Answer #1

Step 1

Moles of I^- present in 20 mL solution = MV

0.08 0,02 1.6 * 103moil

Mole of Ag^+ in 6 mL = 0.1 * 0.006 = 0.6 * 10^-^3 mol

AgI get precipitated and the remaining mol of I^- = (1.6 * 10^-^3) - (0.6 * 10^-^3) = 1.0 * 10^-^3 mol

K_s_p = [Ag^+ ] [I^-]

8.3 * 10^-^1^7 = [Ag^+ ] (10^-^3)

g8.3 10-14

Pag = - log [Ag^+ ] = - log (8.3 * 10^-^1^4) = 13.08

Thus, Pag value is 13.08

Moles of Cl^- present in 20 mL solution = MV

= 0.08 * 0.02 = 1.6 * 10^-^3 mol

Mole of Ag^+ in 6 mL = 0.1 * 0.006 = 0.6 * 10^-^3 mol

AgCI get precipitated and the remaining mol of CI^- = (1.6 * 10^-^3) - (0.6 * 10^-^3) = 1.0 * 10^-^3 mol

K_s_p = [Ag^+ ] [CI^-]

1.77 * 10^-^1^0 = [Ag^+ ] (10^-^3)

[Ag^+ ] = 1.77 * 10^-^7

Pag = - log [Ag^+ ] = - log (1.77 * 10^-^7) = 6.74

Thus, Pag value is 6.74

Step 2

Mole of Ag^+ in 16 mL =0.1 * 0.016 = 1.6 * 10^-^3 mol

AgI get precipitated and the remaining mol of I^- = (1.6 * 10^-^3) - (1.6 * 10^-^3) = 0.0 mol

[Ag^+ ] = [I^-] K_s_p = [Ag^+ ] [I^-]

8.3 * 10^-^1^7 = [Ag^+ ]^2

[Ag^+ ] = 0.93 * 10^-^8

Pag = - log [Ag^+ ] = - log (0.93 * 10^-^8) = 8.03

Thus, Pag value is 8.03

Mole of Ag^+ in 16 mL = 0.1 * 0.016 = 1.6 * 10^-^3 mol

AgCI get precipitated and the remaining mol of I^- = (1.6 * 10^-^3) - (1.6 * 10^-^3) = 0.0 mol

[Ag^+ ] = [CI^-]

K_s_p = [Ag^+ ] [CI^-]

1.77 * 10^-^1^0 = [Ag^+ ]^2

[Ag^+ ] = 1.33 * 10^-^5

Pag = - log [Ag^+ ] = - log (1.33 * 10^-^5) = 4.86

Thus, Pag value is 4.86

Step 3

Mole of Ag^+ in 26 mL = 0.1 * 0.026 = 2.6 * 10^-^3 mole

AgI get precipitated and the extra mol of Ag^+ = (2.6 * 10^-^3) - (1.6 * 10^-^3) = 10^-^3 mol

[Ag^+ ] = M /V

= (10^-^3) /(0.02 + 0.026)

= 2.173 * 10^-^2

Pag = - log [Ag^+ ] = - log (2.173 * 10^-^2) = 1.68

Thus, Pag value is 1.68

Mole of Ag^+ in 26 mL = 0.1 * 0.026 = 2.6 * 10^-^3 mole

AgCI get precipitated and the extra mol of Ag^+ = (2.6 * 10^-^3) - (1.6 * 10^-^3) = 10^-^3 mol

[Ag^+ ] = M /V = (10^-^3) /(0.02 + 0.026) = 2.173 * 10^-^2

Pag = - log [Ag^+ ] = - log (2.173 * 10^-^2) = 1.68

Thus, Pag value is 1.68

Titration curve can be drawn using volume of AgNO_3 used for titration Vs Pag value obtained.